x -2 -1 0 1 2
y .111 .333 1 3 9
Given the table above, graph the function, identify the graph of the function (line, parabola, hyperbola, or exponential), explain your choice, and give the domain and range as shown in the graph, and also the domain and range of the entire function. show work

Answer by scott8148(6628)

the points are (-2,1/9), (-1,1/3), (0,1), (1,3), (2,9) ... 3^-2=1/9, 3^-1=1/3,3^0=1, 3^1=3, 3^2=9

the function is exponential ... y=3^x

the domain of the graph is -2<=x<=2...the range of the graph is (1/9)<=y<=9

the domain of the function is all real numbers ... the range of the function is all real numbers>0